OpenStudy (anonymous):

how to solve (3x-5)/(1-2x)?

OpenStudy (anonymous):

what u meant with solve? u meant find the zeros?

OpenStudy (anonymous):

(3x-5)/(1-2x)=0? then (3x-5)=0*(1-2x) multiply both sides by (1-2x) 3x-5=0 3x=5 x=5/3